# Prediction of the Most Fire-Sensitive Point in Building Structures with Differentiable Agents for Thermal Simulators

### v1

@@ IV-B Material Properties
IV-C Gravity Loads
Gravity loads are applied to columns and beams based on their position and function. For columns, only self-weight is considered, while beams support additional loads from reinforced concrete slabs, people as live load, and building content. Edge beams, which support only one side, carry less load than interior beams. The ranges of gravity loads are listed in Table II. Structures that failed to meet a MIDR threshold of $1\%$ under gravity loads were deemed unstable and filtered out, as such configurations of randomly chosen geometry, material, and gravity load combinations were considered unrealistic from a regulatory and practicality points of view.
TABLE II: Gravity load ranges for beams and columns.
@@ IV-C Gravity Loads
IV-D Rule-based Thermal Load Generation
To evaluate a building’s structural response during a fire, we employed a simplified rule-based approach to thermal load generation. Previous studies [23] have demonstrated that steel structures rapidly equilibrate with surrounding gases temperatures due to efficient heat exchange. Consequently, gas temperatures can be directly used as inputs for FEA tools, e.g., OpenSees, simplifying the process of modeling thermal loads.
Accurately simulating temperature fields in fire scenarios poses significant challenges. Advanced thermodynamic simulations, such as those performed using Fire Dynamics Simulator (FDS) [20], provide precise temperature predictions. However, these methods are hindered by high computational costs, prolonging execution times, and limited scalability, making them impractical for generating large datasets. Additionally, real-world fire loads often display substantial spatial variability across different rooms [7], resulting in scenario-specific temperature fields with limited generalizability. For example, studies on bridge fires [11] have demonstrated that environmental factors, such as wind speeds, can significantly influence temperature distributions. Furthermore, even within identical scenarios, variations in fire modeling methodologies can produce distinctly different temperature fields [27, 6]. These challenges emphasize the need for efficient and adaptable methods to generate fire temperature data.
To address these issues, we adopted a rule-based approach to model temperature variations. According to [25], a typical fire development follows a predictable pattern: during the growth stage, the temperature rises slowly and approximately linearly after ignition. This is followed by the flashover stage, where temperatures increase rapidly to peak values. After reaching the peak, the temperature either stabilizes or continues to rise slowly until the decay stage begins. Inspired by this fire development pattern, we describe the temperature evolution prior to decay in two distinct stages:
@@ IV-D Rule-based Thermal Load Generation
The slope of the linear temperature increase, $c$, and the transition time, $t_{1}$, are influenced by the spatial relationship between the fire source and the structural element. For the second stage of temperature evolution, we utilize the ISO 834 curve, a widely accepted standard for fire resistance testing. This standardized fire curve describes the temperature rise over time, enabling rapid and consistent thermal fields across various scenarios. Additionally, the duration of fire simulation in this study is set to $t_{\text{th}}=60$ minutes. This value represents the upper time limit for the temperature evolution of each structural element, providing a consistent basis for analyzing structural responses to fire.
Let $(x,y,z)$ represents the midpoint of a structural element and $(x_{\text{f}},y_{\text{f}},z_{\text{f}})$ the fire origin point. Heights $h$ and $h_{\text{f}}$ correspond to the respective floor levels of the element and the fire source. The temperature evolution for each element is expressed as:
@@
The transition (arrival) time $t_{1}$, marking the end of the linear stage, depends on the spatial distance between the fire source and the element. We define the following two Euclidean distances $L_{p}$ in the $xy$ plane and $L_{s}$ in the $xyz$ space:
@@
The parameters $\beta$ and $\alpha$ for determining $t_{1}$ are summarized in Table III. In this study, we take $r_{\mathrm{up}}=0.95$ and $r_{\mathrm{down}}=0.97$.
@@
Figure 4 illustrates the $t_{1}$ curves for various fire scenarios: fire originating on the lower floor, $h-h_{\text{f}}=1$ with rapid upward spread, fire on the same floor, $h=h_{\text{f}}$ with the fastest spread, and fire on the upper floor, $h_{\text{f}}-h=1$ with slower downward propagation. The exponential decay observed in $t_{1}$ reflects the accelerating fire propagation speed as the distance increases. Figure 4 also indicates that the employed simplified model is consistent with the Markov chain-based dynamic model given by [4], where the rooms at the same floor of the fire point start flashover slightly before the corresponding upper floors. Additionally, $\beta_{1}$ and $\beta_{3}$ are the summation of a geometric sequence, where story level $h$ is the index. The common ratios $r_{\mathrm{up}}<1$ in $\beta_{1}$ and $r_{\mathrm{down}}<1$ in $\beta_{3}$ indicate that the fire speeds up to spread through the next story, which is consistent with the real-world fire spread mechanism given in [12]. The temperature profile within the range $t\in[0,t_{\text{th}}]$ is subsequently used as the thermal load in OpenSeesRT simulations to compute displacements at each structural node at time $t_{\text{th}}$.
@@
In conclusion, this rule-based approach is a computationally efficient method to approximate fire temperature fields, enabling large-scale dataset generation to train predictive models. By combining ISO 834 fire curves with spatial considerations, the method balances accuracy and scalability, making it a practical solution for thermal load modeling in fire scenarios.
IV-E Interstory Drift Ratio
The Interstory drift ratio (IDR) quantifies the relative displacement between two consecutive floors (interstory displacement) as a percentage of the floor height. It serves as a critical parameter for evaluating structural deformation under external forces, such as seismic or thermal loads. In the context of fire-induced deformation, the IDR provides valuable insight into the relative movement between floors as a structure experiences thermal expansion or contraction. Mathematically, the IDR for a given node $i$, i.e., $d_{i}$, is defined as follows:
where the numerator represents the relative displacement (in the horizontal plane $xy$) of node $i$ with respect to the corresponding node on the floor below, and $H$ denotes the story height between these two floors. The IDR is a crucial metric for assessing the structural integrity under extreme conditions. Excessively high drift ratios can indicate significant structural deformation, potentially leading to damage or collapse. For this study, the IDR at $t_{\text{th}}$ is computed using displacement data obtained from OpenSees simulations under thermal loading during fire scenarios.
IV-F OpenSees Simulation
The thermal and mechanical responses of 3D frame structures under combined fire and gravity loads are simulated using OpenSees [21], specifically OpenSeesRT [24]. Each structural model features six degrees of freedom per node, with linear geometric transformations (geomTransf: Linear) defining element orientations. Material behavior is temperature-dependent and modeled with Steel01Thermal, while fiber-based sections (FiberThermal) capture nonlinear interactions between thermal and mechanical responses at the cross-section level. Structural elements are represented as displacement-based beam-column elements (dispBeamColumnThermal), accounting for both thermal and mechanical load effects.
@@ IV-F OpenSees Simulation
The analysis begins with the application of gravity loads to establish equilibrium, followed by incremental thermal loads simulating fire exposure. A static nonlinear solver using the ExpressNewton algorithm ensures efficient convergence, while the NormDispIncr test maintains accuracy. An incremental LoadControl scheme with small step sizes is employed to guarantee numerical stability, using 10% for gravity loads and 1% for thermal loads. In the thermal load analysis of steel structures using OpenSeesRT, the material properties (e.g., Young’s modulus and yield strength) are adjusted at increasing temperatures according to [8], particularly its Table 3.1: Reduction factors for the stress-strain relationship of carbon steel at elevated temperatures. OpenSees documentation [9] provides several examples of thermal analysis for different steel materials.
@@
This modeling framework accommodates variations in material properties, cross-sectional geometries, and temperature profiles, providing robust simulations of structural behavior under fire conditions. The primary settings and configurations for the OpenSeesRT simulations are summarized in Table IV.
TABLE IV: Key settings of OpenSeesRT simulation scripts.
@@
Geometry 3D models; 6 degrees of freedom per node
@@
Section FiberThermal; Cross-section dimensions: $0.1$ m $\times$ $0.1$ m
@@
For each structure in the labeled dataset, 30 fire points are selected using a dual-granularity approach to ensure they are well-distributed. Specifically, rooms are sequentially selected, with one fire point randomly chosen within each selected room. If a building is large and contains more than 30 rooms, we randomly select 30 rooms without replacement, i.e., ensuring that no more than one fire point is located in the same room. Conversely, if the building is small and has fewer than 30 rooms, all rooms are initially selected, with one fire point randomly assigned to each room. Additionally, rooms are then selected with replacement until a total of 30 fire points are assigned. This selection method prevents fire points from clustering too closely while maintaining an element of randomness. By distributing fire points in this manner, the 30 fire scenarios are effectively utilized, enabling a more comprehensive characterizing of how the structures respond under fire conditions.
IV-G Summary of the Dataset Generation
As discussed in this section and related to Figure 2, three key steps were considered in the development of the dataset:
@@ IV-G Summary of the Dataset Generation
    Filtering Process: Structures with MIDR exceeding $1\%$ under gravity loads were excluded, resulting in $1,573$ labeled structures retained for fire simulation and $16,050$ unlabeled structures for training the MFSP Predictor.
@@
    Fire Simulations: For each retained labeled structure, 30 fire scenarios were simulated using OpenSeesRT, yielding $47,190$ fire cases.
@@
    Data Distribution Check: MIDR distributions for labeled and unlabeled data under gravity loads were highly similar, because both datasets were generated using the same method. Under fire conditions, the MIDR distribution shifted, reflecting significant structural deformation with values reaching a maximum of about 6%, an average of 1.70%, and a standard deviation of 1.12%. This step ensured a diverse and comprehensive dataset for the predictive framework.
@@
The statistical distribution histograms for MIDR (after applying the $1\%$ filtering threshold) under different loading conditions are plotted in Figure 5 to illustrate the data characteristics. Figures 5(a) and 5(b) show the MIDR distributions of the labeled and unlabeled data, respectively, under gravitational loads only. Figure 5(c) depicts the MIDR distribution of the labeled data under the combined effects of gravity and fire loads. Under fire loading, the structures exhibit significant deformation, leading to a noticeably right-shifted MIDR distribution.

### v5

@@ VI-B Material Properties
VI-C Gravity Loads
Gravity loads are applied to columns and beams based on their influence (tributary) areas as typically conducted in structural design. The considered load conditions include the column self-weight and the additional loads directly supported on the beams from their self-weight and weights of the reinforced concrete slabs, and building content. An edge beam typically carries approximately half the gravity load supported by a parallel interior beam. The gravity loads are set to be uniformly sampled from ranges to reflect real-world variability, as listed in Table II. Structures that failed to meet an MIDR threshold of $1\%$ under gravity loads are deemed unacceptable designs and filtered out, as such configurations of randomly chosen geometry, material, and gravity load combinations are considered unrealistic from a regulatory and practicality points of view.
TABLE II: Gravity load ranges for considered beams and columns.
@@ VI-C Gravity Loads
VI-D Rule-based Thermal Load Generation
To evaluate a building’s structural response during a fire event, we employ a simplified rule-based approach for thermal load generation. According to [38], a typical fire development follows a predictable pattern. During the growth stage after ignition, the temperature rises slowly and approximately linearly. This is followed by the burning stage starting from the onset of flashover, where temperatures increase rapidly to peak values. After reaching the peak, the temperature either stabilizes or continues to rise slowly until the decay stage begins. Inspired by this fire development pattern, we describe the temperature evolution in time, $t$, prior to the decay stage in two distinct stages:
@@ VI-D Rule-based Thermal Load Generation
The slope of the linear temperature increase, $c$, and the transition time, $t_{1}$, are influenced by the spatial relationship between the fire source and the structural element. For the second stage of temperature evolution, we utilize the ISO 834 curve, a widely accepted standard for fire resistance testing. This standardized fire curve describes the temperature rise over time, enabling rapid and consistent thermal fields across various scenarios. The duration of fire simulation in this study is set to $t_{\text{th}}=60$ minutes. This value represents the upper limit for the temperature evolution of each structural element, providing a consistent basis for analyzing the structural response to fire.
Let $(x,y,z)$ represents the midpoint of a structural element and $(x_{\text{f}},y_{\text{f}},z_{\text{f}})$ the fire source point. Integer parameters $h$ and $h_{\text{f}}$ correspond to the respective floor levels of the element and the fire source. The temperature evolution for each element is expressed as follows:
@@
The transition (arrival) time $t_{1}$ in minutes, i.e., end of the linear stage, depends on the spatial distance between the fire source and the element. We define the following two Euclidean distances $L_{p}$ in the $X-Y$ plane and $L_{s}$ in the $XYZ$ space:
@@
The parameters $\beta_{i}$ (same time unit as $t_{1}$) and $\alpha_{i}$ (same length unit as $L_{s}$ or $L_{p}$) for determining $t_{1}$ are summarized in Table III. In this study, we take $r_{\mathrm{up}}=0.95$ and $r_{\mathrm{down}}=0.97$.
@@
Figure 7 illustrates the $t_{1}$ curves ($t_{1}$ in minutes vs. $L_{s}$ or $L_{p}$ in meters) for various fire scenarios: (1) fire originating on the lower floor, $h-h_{\text{f}}=1$ with rapid upward spread, (2) fire on the same floor, $h=h_{\text{f}}$ with the fastest spread, and (3) fire on the upper floor, $h_{\text{f}}-h=1$ with slow downward spread. The exponential decay terms in $t_{1}$ relationships (Equations 18) capture the acceleration of fire spread as the distance increases. As shown in Figure 7, the adopted simplified model aligns well with the dynamic model based on Markov chains presented in [7], wherein rooms on the same floor as the ignition point undergo flashover slightly earlier than those on the floors above. Moreover, both $\beta_{1}$ and $\beta_{3}$ can be expressed as the sum of geometric series, with the story level $h$ serving as the index. The fact that the ratios $r_{\mathrm{up}}<1$ in $\beta_{1}$ and $r_{\mathrm{down}}<1$ in $\beta_{3}$ implies an increasing rate of vertical fire spread across stories, which is consistent with the empirical fire propagation mechanisms discussed in [19]. The temperature profile within the range $t\in[0,t_{\text{th}}]$ is subsequently used as the thermal load in $\chi$ara simulations to compute the displacements at each structural node at time $t_{\text{th}}$.
@@
Remark: The effects of structural elements, such as concrete floor slabs and partitions, are not explicitly modeled in our approach. Instead, their influence is implicitly captured through the careful selection of the parameters $\alpha_{i},\beta_{i},r_{\mathrm{up}}$, and $r_{\mathrm{down}}$. This parameterization provides a unified framework for generating temperature fields. Indeed, fire propagation is governed by a multitude of factors and remains an open research question. For instance, if the fire resistance of a floor slab is enhanced by fire protective coating, the corresponding model can account for this by decreasing $\alpha_{1}$ $\alpha_{3}$, increasing $\beta_{1}$ $\beta_{3}$, and adopting larger values for $r_{\mathrm{up}}$ $r_{\mathrm{down}}$, which collectively slows down the vertical spread of fire. Conversely, scenarios involving higher amounts of combustible materials would warrant the opposite adjustments. This flexible and integrated approach avoids the need to design and analyze separate models for different fire propagation scenarios while still capturing the essential effects.
@@
In conclusion, our rule-based approach is a computationally efficient method for approximating fire temperature fields, enabling large-scale dataset generation to train predictive models. By combining ISO 834 fire curves with spatial considerations and embedding structural effects through parameter calibration, the method achieves a balanced trade-off between accuracy and scalability, making it a practical solution for thermal load modeling in fire scenarios. After generating the temperatures of the midpoints of each beam and column, these temperatures are applied as uniform thermal load to these elements of the building structure in question using $\chi$ara.
VI-E $\chi$ara Simulation
The thermal and mechanical responses of 3D frame structures under combined fire and gravity loads are simulated using $\chi$ara In the simulation, the IDR of each node at $t_{\text{th}}$ is determined using the computed nodal displacements. Each structural model features 6 Degrees of Freedom (DoFs) per node (3 translations & 3 rotations), with linear geometrical transformations (geomTransf: Linear) defining how the element local coordinate systems are mapped to the global coordinate system and assuming small displacements and rotations. Although $\chi$ara allows a variety of options for modeling finite deformations, in the present simulations and mainly for simplicity, we did not consider large deformations. All bottom nodes (on the ground) are fully constrained in all 6 DoFs, while those of all other nodes are free. Material behavior is temperature-dependent and modeled with Steel01Thermal, while fiber-based sections (FiberThermal) capture nonlinear interactions between thermal and mechanical responses at the cross-section level. Structural elements are modeled as displacement-based Euler-Bernoulli beam-columns (dispBeamColumnThermal). This element formulation accounts for thermal strains (temperature gradients) in the section, which is discretized into fibers. Numerical integration is used along the length of each element using three integration (Gauss) points, one at each end and the third in the middle of the element.
Thermal expansion of steel members plays a crucial role in IDR development. In reality, reinforced concrete floor slabs heat at a different rate than steel members due to their higher thermal mass and lower thermal conductivity. This differential heating can lead to restrained thermal expansion, introducing axial compression in beams and affecting the overall structural response. In this study, explicit composite action between steel members and concrete slabs is not modeled. Instead, our approach focuses on isolating the response of the steel structural frame, which is often the critical load-bearing component in fire scenarios. This assumption aligns with prior studies [34] demonstrating that steel structures reach thermal equilibrium with surrounding gases quickly, allowing the use of uniform thermal loading in fire analysis. Future work could enhance this framework by incorporating slab-beam interaction effects, through a refined FEA for an extended dataset where constraints imposed by floor slabs are explicitly considered.
@@ VI-E $\chi$ara Simulation
The analysis begins with the application of gravity loads, followed by incremental thermal loads simulating the fire exposure. A static nonlinear solver using ExpressNewton algorithm ensures convergence, while the NormDispIncr test maintains accuracy. An incremental LoadControl scheme with small step sizes is employed to guarantee numerical stability, using 10% for gravity loads and 1% for thermal loads.
In the thermal load analysis, uniform thermal load is applied to each beam and column, i.e., the temperature of each element is set to be that at the middle point, according to Section VI-D. The Steel01Thermal material allows the properties (e.g., Young’s modulus and yield strength) to be adjusted at increasing temperatures according to [11] using its Table 3.1: Reduction factors for the stress-strain relationship of carbon steel at elevated temperatures. For example, if Young’s modulus at ambient temperature is $E_{0}$, then as the temperature $T$ increases, the modulus changes as $E(T)=\eta(T)\times E_{0}$. Reference [11] directly provides the values of $\eta(T)\in\left[0,1\right]$ at every $100^{\circ}\mathrm{C}$ interval and recommends linear interpolation to obtain $\eta(T)$ for intermediate $T$ values. OpenSees documentation [12] provides several examples of thermal analyses.
@@
This modeling framework accommodates variations in material properties, cross-sectional geometries, and temperature profiles, providing robust simulations of structural behavior under fire conditions. The primary settings and configurations for the $\chi$ara simulations are summarized in Table IV.
TABLE IV: Key settings of $\chi$ara simulations.
@@
Geometry 3D models; 6 DoFs per node
@@
Section FiberThermal; Cross-section: $0.1$ m $\times$ $0.1$ m
@@
For each structure in the labeled dataset, 30 fire points are selected using a dual-granularity approach, i.e., two-stage sampling strategy, to ensure they are well-distributed. Specifically, rooms are sequentially selected, with one fire point randomly chosen within each selected room. If a building is large and contains more than 30 rooms, we randomly select 30 rooms without replacement, i.e., ensuring that no more than one fire point is located in the same room. Conversely, if the building is small and has fewer than 30 rooms, all rooms are initially selected, with one fire point randomly assigned to each room. Additionally, rooms are then selected with replacement until a total of 30 fire points are assigned. The room-level sampling prioritizes selecting distinct rooms to avoid spatial clustering of fire points, while the point-level sampling ensures intra-room variability. This approach aligns with stratified sampling principles commonly used for efficient spatial representation, where multi-stage sampling strategies optimize coverage and variability, e.g., [4], and enables a more comprehensive characterization of how the structures respond under fire conditions.
VI-F Summary of the Dataset Generation
As discussed in this section and related to Figure 5, three key steps are considered in the development of the dataset:
@@ VI-F Summary of the Dataset Generation
    Filtering process: Structures with MIDR exceeding $1\%$ under gravity loads are excluded, resulting in $1,573$ labeled structures retained for fire simulations and $16,050$ unlabeled structures for training the MFSP predictor.
@@
    Fire simulations: For each retained labeled structure, 30 fire scenarios are simulated using $\chi$ara ($47,190$ fire cases).
@@
    Data distribution check: MIDR distributions for labeled and unlabeled data under gravity loads are highly similar because both datasets are generated using the same method. Under fire conditions, the MIDR distribution shifted, reflecting significant structural deformation with values reaching a maximum of about 6%, an average of 1.70%, and a standard deviation of 1.12%. This step ensured a diverse and comprehensive dataset for the proposed predictive framework.
@@
The statistical distribution histograms for MIDR (after applying the $1\%$ filtering threshold for gravity load responses) under different loading conditions are plotted in Figure 8. Figures 8(a) and 8(b) show the MIDR distributions of the labeled and unlabeled data, respectively, under gravity loads only. Figure 8(c) shows the MIDR distribution of the labeled data under the combined effects of gravity and fire loads. Clearly, fire loads cause the structures to significantly deform, leading to a noticeably right-skewed MIDR distribution.